OpenStudy (anonymous):

Solve the following system of equations graphically. Click on the graph until the correct one appears. 2x -y = -1 x + y = -2

OpenStudy (anonymous):

For the following system, use the second equation to make a substitution for y in the first equation. 3x + y = 2 y + 4 = 3x

OpenStudy (anonymous):

same method ... try it ..

OpenStudy (anonymous):

Given the system of equations, what is the solution? 2x + y = -1 x - y = -5

OpenStudy (anonymous):

Check out this graph

OpenStudy (anonymous):

The red one represents 2x -y = -1 The blue line represents x + y = -2

OpenStudy (anonymous):

One number is five more than three times another. If their difference is 35, what is the larger number?

OpenStudy (anonymous):

https://godleyisd.owschools.com/owsoo/studentAssignment/index unit 13 final exam algebra 1

OpenStudy (anonymous):

Take one number as x and the other as y. x=3y+3 x-y=35 Solve both the equations and you can get the answer y=16 x=51

OpenStudy (anonymous):

Man I deserve a medal for this hardwork